# Error Crafting in Probabilistic Quantum Gate Synthesis

### v1

@@ Acknowledgement
It can be shown that this relates a single-qubit unitary channel $\mathcal{U}$ to a unit vector $r\in\mathbb{R}^{4}$ as follows,
@@
S2 Numerical details on error crafting using unitaries
S2.1 Crafting Pauli channel
As discussed in the main text, the minimization problem regarding error-crafted synthesis of a Pauli channel is formulated via the following constrained minimization problem of a probability distribution $\{p_{j}\}$ given a target unitary channel $\hat{\mathcal{U}}$ and a set of synthesized unitaries $\{\mathcal{U}_{j}\}$ as follows:
@@
When the set $\{\mathcal{U}_{j}\}$ is not appropriately generated, there does not exist any feasible solution for Eq. (S19). While numerical solvers such as Gurobi offers an automated way to relax the equation constraint, via the variable presolve, we find that this precomputation is done too aggressively so that the accuracy of the minimization is quite poor. Instead, in our work, we have done a logarithmic search of $g_{1}$, starting from machine precision of $g_{1}=10^{-16}$. We have defined the error crafting to be successful if we find feasible solution such that (1) $\min\frac{1}{2}\|\hat{\mathcal{U}}-\sum_{j}p_{j}\mathcal{U}_{j}\|_{\diamond}\leq(c+1)^{2}\epsilon^{2},$ and (2) $g_{1}\leq 10^{-12}$ are both satisfied.
We remark that, using the magic basis introduced in Eq. (S18), the minimization problem boils down to a linear programming problem as
@@
where $r^{(j)}=(r_{1}^{(j)},r_{2}^{(j)},r_{3}^{(j)},r_{4}^{(j)})^{T}\in\mathbb{R}^{4}$ corresponds to the magic basis representation of $\mathcal{U}_{j}\circ\hat{\mathcal{U}}^{-1}$. We find that accuracy is much enhanced when the minimization (and relaxation) is done under this representation. In practice, we solve Eq. (S21) using LP implemented in Gurobi, and check if the conditions are satisfied.
S2.2 Crafting depolarizing channel
Here, we provide numerical details regarding the error crafting of depolarizing channel. In similar to Sec. S6.1, the minimization problem is defined as
@@
${\rm minimize}\ \frac{1}{2}\left\|\hat{\mathcal{U}}-\sum_{j}p_{j}\mathcal{U}_{j}\right\|_{\diamond}\ \ \ {\rm subject\ to}\ \ \sum_{a}\chi_{a}=1,\max_{a^{\prime}\neq b^{\prime}}\{|\chi_{a^{\prime}a^{\prime}}-\chi_{b^{\prime}b^{\prime}}|\}\leq g_{2},\sum_{a\neq b}|\chi_{a\neq b}|\leq g_{1},$ (S23)
@@
where $a\in\{{\rm I},{\rm X},{\rm Y},{\rm Z}\}$ and $a^{\prime},b^{\prime}\in\{{\rm X},{\rm Y},{\rm Z}\}.$ We regard the error crafting to be successful if (1) $\min\frac{1}{2}\|\hat{\mathcal{U}}-\sum_{j}p_{j}\mathcal{U}_{j}\|_{\diamond}\leq(c+1)^{2}\epsilon^{2}$, (2) $g_{1}\leq 10^{-12}$, and (3) $g_{2}\leq 0.01(\sum_{a^{\prime}\in\{X,Y,Z\}}\chi_{a^{\prime}a^{\prime}})\sim 0.01\epsilon^{2}$ are all satisfied.
@@
The results of error-crafted synthesis is shown in Fig. S1. We observe the validity of the nine shift unitary from the bounded behavior in Fig. S1(a). As in the case for Pauli constraint, we also observe a boost in the accuracy by expanding the size of the basis set by factor of $R$. From Figs. S1(b), (c), and (d), we further find that both increasing $c$ and $R$ contributes to suppress the failure probability of the error crafting.
@@
Figure S1: (a) Diamond distance $d_{\diamond}(\hat{\mathcal{U}},\mathcal{U}_{\rm synth})$ between the target single-qubit Haar random unitary $\hat{\mathcal{U}}$ and synthesized channel $\mathcal{U}_{\rm synth}\coloneqq\sum_{j}p_{j}\mathcal{U}_{j}$, with the remnant error constrained to the single-qubit depolarizing channel. The black real line indicates the results from unitary synthesis by the Ross-Selinger algorithm [42], while the filled dots are results from error-crafted synthesis with shift factors of $c=2,3,5,7.$ When we expand the basis set size by factor of $R=3$, shown by unfilled triangles, the effective synthesis error is suppressed below the bound in Eq. (S58) and approaches the value of $\epsilon^{2}$. The shown data are mean over 200 trials. (b) Failure rate of error-crafted synthesis, averaged over 50 trials. (c)(d) Suppression of $p_{\rm fail}$ with shift factor $c$ and $R$. The data is averaged over 10000 and 2000 trials, respectively.
@@
S3 Numerical details on error crafting using CPTP maps
As discussed in the main text, the error of Pauli rotation can be suppressed significantly by applying the error crafting for CPTP maps that represents the channel of synthesized gates accompanied by error correction circuit as shown in Fig. S2. When the target unitary and synthesized unitary channels are given as $\hat{\mathcal{U}}$ and $\mathcal{U}_{j}$, respectively, the effective CPTP channel $\Lambda_{j}$ and remnant error $\mathcal{E}_{{\rm rem},j}$ obtained from logical error correction can be written as
@@
In order to clarify how to perform the error crafting for a single-qubit Pauli Z rotation, first note that the Pauli transfer matrix of a remnant error channel $\mathcal{E}_{{\rm rem},j}$ for a Pauli Z rotation, denoted as $\Gamma_{\mathcal{E}_{{\rm rem},j}}$, can be expressed as
@@
where $0\leq\mu_{j}\leq 1$ and $\Gamma_{\mathcal{E}_{{\rm rem},j}}\geq 0.$ This motivates us to find a pair of remnant error channels $\mathcal{E}_{{\rm rem},+}$ and $\mathcal{E}_{{\rm rem},-}$ which satisfies (i) $\nu_{+}>0,\nu_{-}<0$, (ii) $\mu_{\pm}=O(\epsilon^{3})$. Once such a pair is found, we take the probabilistic mixture of them as
@@
which is a purely Z error channel with error rate of $\tilde{\mu}=\frac{\nu_{-}}{\nu_{+}+\nu_{-}}\mu_{+}+\frac{\nu_{+}}{\nu_{+}+\nu_{-}}\mu_{-}=O(\epsilon^{3}).$
We note that the current search algorithm is based on a brute-force search, and hence the computational complexity is expected to scale as $O({\rm poly}(1/\epsilon))$, while we leave an open question whether there is any bound on complexity to search the pair as in the above discussion. In the current implementation, we generate around 3000 unitaries to reach remnant error of $10^{-6}$ and around 20000 unitaries to reach $10^{-9},$ with each unitary synthesis done in milliseconds using a direct search algorithm proposed in Ref. [51]. Note that each unitary synthesis is independent, and hence is embarrassingly parallelizable.
S4 Generation of shift unitaries with fixed diamond distance
@@
S5 White noise approximation under coherent errors
@@
The white noise has been argued to be realized in experiments, in particular when the unitary gates are chosen at random [49]. In fact, one can prove that, even if the qubit connectivity is only linear and the circuit constitutes a brick-wall structure with random Haar unitary gates for two qubits, the effective noise converges to the white noise under unital noise [50, 60]. However, these arguments mainly consider incoherent error, and how coherent error can affect the white-noise approximation is not fully examined. Given that the coherent component of error arises as the synthesis error in the early FTQC regime, it is crucial to investigate the effect of coherent errors on the white-noise approximation.
@@
by choosing the rescaling factor as $R=(s/u)^{L}$ (Theorem S1 of [61]). Here, $u$ and $s$ are noise-dependent constants called unitarity and average noise strength, respectively, which can be represented as
@@
We can observe a similar scaling numerically. Before discussing the effect of the coherent noise, let us briefly summarize the numerical argument for incoherent noise, especially the local depolarizing noise $\mathcal{E}=\bigcirc_{i=1}^{n}\mathcal{E}^{(i)}_{\mathrm{dep}}$, provided in Ref. [33]. We analyze the convergence to white noise by examining the Pauli transfer matrix (PTM) of the effective noise $\mathcal{E}^{\prime}$, denoted hereafter as $\Gamma$. We define $\Gamma$ by $\Gamma_{ij}={\rm Tr}[P_{i}\mathcal{E}^{\prime}(P_{j})]/2^{n}$ for $P_{i},P_{j}\in\{I,X,Y,Z\}^{\otimes n}$ for $n$-qubit system. Note that PTM is diagonal if $\mathcal{E}^{\prime}$ is a Pauli channel, and furthermore there exists $0\leq w\leq 1$ such that $\Gamma_{ii}=w~{}(i\neq 0)$ if $\mathcal{E}^{\prime}$ is the white noise. In this regard, it is useful to compute the singular values of the PTM in order to investigate the closeness to the white noise. We in particular compute $\Gamma=V\Lambda W^{\dagger}$ with $\Lambda={\rm diag}\{\lambda_{i}\}_{i=1}^{4^{n}}$, and define the damping factor $k_{i}$ as follows,
@@
It can be proven that, when each unitary layer consists of unitary 2-design, the damping factor converges to $k_{\rm mean}=\frac{3n}{4}\frac{4^{n}}{4^{n}-1}$, and numerical analysis supports that this also holds for weaker randomness, e.g., 2-local random circuits with linear connectivity, when the circuit depth is sufficiently deep. In other words, the mean deviation $\mathbb{E}_{i}[|k_{i}-k_{\rm mean}|]$, as well as the deviation of extremals $|k_{\rm max(min)}-k_{\rm mean}|$, is suppressed as $O(1/\sqrt{L})$ [33], which is consistent with the suppression of total variation distance in measurement probability distribution in the computational basis [50].
Now we ask how the above behavior is affected by the presense of coherent errors. Here, for the sake of simplicity of the argument, we sample $U_{i}$ uniformly from the entire unitary group, i.e., take an $n$-qubit Haar random unitary, and consider a local noise as $\mathcal{E}_{l}=\bigcirc_{i=1}^{n}\mathcal{E}_{\rm coh}^{(i)}\circ\mathcal{E}_{\rm dep}^{(i)},$ where $\mathcal{E}_{\rm coh}^{(i)}$ is a random 1-qubit unitary such that $\|\mathcal{E}_{\rm coh}^{(i)}-{\rm Id}\|_{\diamond}=\epsilon_{\rm coh}.$ In practice, this can be realized by constraining the rotation angles under axial decomposition of the unitary.
Interestingly, we find that the effect of coherent error on the magnitude of the white noise is quadratically suppressed, while that on the convergence rate is linear. In Fig. S3(a)(b) we can see that, although the fluctuation in the damping factor is expanded at the shallow-depth regime, it readily converges at the rate of $O(1/\sqrt{L})$, which is in similar to the depolarizing-only case [33]. We observe that the deviation $\max_{k}\{|k-k_{\rm mean,dep}|\}$ is lifted approximately by a factor of $(\epsilon_{\rm coh}+\epsilon_{\rm dep})/\epsilon_{\rm dep}$ where $\epsilon_{\rm dep}:=\|\mathcal{E}_{\rm dep}^{(i)}-{\rm Id}\|_{\diamond}$, implying the slow down of the white noise convergence. Meanwhile, as shown in Fig. S3(c), we notice that the converged effective channel is less affected, in a sense that $|k_{\rm mean,tot}-k_{\rm mean,dep}|\propto\epsilon^{2}.$ Phenomenologically, we can understand that the effect of coherent noise is self-twirled by the circuit such that the diamond norm is suppressed quadratically as in the random compilation scheme [62]. These results indicate that, although the presence of coherent error does not severely increase the sampling overhead to perform error countermeasures such as the rescaling technique, the adequacy of the approximation is degraded and hence desirable to avoid them by error-crafted synthesis.
@@
S6 Theoretical guarantees for error crafted synthesis
S6.1 Crafting Pauli channel (Proof of Theorem 1)
@@
minimize ${\rm minimize}\ 1-\sum_{j}p_{j}\left(r^{(j)}_{1}\right)^{2}\ \ \ {\rm subject\ to}\ \ \sum_{j}p_{j}r^{(j)}r^{(j)T}\ {\rm is\ diagonal},$ (S36)
@@
where $\hat{A}_{6}$ is the matrix defined in the same way as Eq. (S39) for $\hat{\mathcal{U}}_{j}\circ\hat{\mathcal{U}}^{-1}$, and we use $\hat{A}_{6}^{-1}\hat{m}^{(7)}=(D\hat{A}_{6})^{-1}D\hat{m}^{(7)}$ in the last equality. If we can show that there exists an upper bound $u(c,\epsilon)$ of $\left\|Dm^{(j)}-D\hat{m}^{(j)}\right\|_{2}$ satisfying
@@
where we use the Cauchy-Schwartz inequality $\left(\sum_{i=1}^{3}\left(\hat{v}_{i}^{2}-\frac{\sin^{2}\theta}{\sin^{2}\hat{\theta}}v_{i}^{2}\right)^{2}\right)\left(\sum_{i=1}^{3}\left(\frac{1}{\sqrt{3}}\right)^{2}\right)\geq\left(\sum_{i=1}^{3}\frac{1}{\sqrt{3}}\left(\hat{v}_{i}^{2}-\frac{\sin^{2}\theta}{\sin^{2}\hat{\theta}}v_{i}^{2}\right)\right)^{2}$ in the last inequality. We proceed the calculation as follows:
@@
S6.2 Crafting depolarizing channel
@@
minimize ${\rm minimize}\ 1-\sum_{j}p_{j}\left(r^{(j)}_{1}\right)^{2}\ \ \ {\rm subject\ to}\ \ \sum_{j}p_{j}r^{(j)}e^{(j)T}=(1-q)\mathbf{e}^{(1)}\mathbf{e}^{(1)T}+\frac{q}{4}{\rm Id},$ (S56)
@@
S7 Technical lemmas
S7.1 Uniform continuity of $g_{c}(\epsilon,t)$
@@
S7.2 Commutativity of $\lim$ and $\max$ for uniformly continuous functions

### v2

@@ Acknowledgement
It can be shown that this relates a single-qubit unitary channel $\mathcal{U}$ to a unit vector $r=(r_{1},r_{2},r_{3},r_{4})^{T}\in\mathbb{R}^{4}$ as follows,
@@
When we explicitly consider the distance between a target unitary $\hat{\mathcal{U}}$ and a mixture of unitaries $\sum_{j}p_{j}\mathcal{U}_{j}$, we can rewrite the expression of the diamond distance as
$d_{\diamond}\left(\hat{\mathcal{U}},\sum_{j}p_{j}\mathcal{U}_{j}\right)=\frac{1}{2}\left\|\hat{\mathcal{U}}-\sum_{j}p_{j}\mathcal{U}_{j}\right\|_{\diamond}=\frac{1}{2}\cdot\frac{1}{2}\left\|\mathcal{J}(\mathcal{U})-\sum_{j}p_{j}\mathcal{J}(\mathcal{U}_{j})\right\|_{1}=\frac{1}{2}\left\|rr^{T}-\sum_{j}p_{j}r^{(j)}r^{(j)T}\right\|_{1},$ (S23)
where $r$ and $r^{(j)}$ are the unit vectors obtained from the magic basis representation of $\hat{\mathcal{U}}$ and $\mathcal{U}_{j}$, respectively. Furthermore, since the trace norm $\|A\|_{1}$ is simply a sum of diagonal elements when $A$ is diagonal, the diamond distance can be simplified when $\hat{\mathcal{U}}$ is identity and $\sum_{j}p_{j}\mathcal{U}_{j}$ is a single-qubit Pauli channel as
$d_{\diamond}\left(\hat{\mathcal{U}},\sum_{j}p_{j}\mathcal{U}_{j}\right)=\frac{1}{2}\left((1-\sum_{j}p_{j}(r_{1}^{(j)})^{2})+\sum_{k=2}^{4}\sum_{j}p_{j}(r_{k}^{(j)})^{2}\right)=1-\sum_{j}p_{j}(r_{1}^{(j)})^{2},$ (S24)
where we have used $\sum_{k=2}^{4}(r_{k}^{(j)})^{2}=1-(r_{1}^{(j)})^{2}$ for the second equation. As we discuss in Sec. S3, this simplification is beneficial since the minimization problem of the diamond distance can be formulated by linear programming problem instead of the cumbersome SDP.
S3 Numerical details on error crafting using unitaries
S3.1 Crafting Pauli channel
As discussed in the main text, the minimization problem regarding the error-crafted synthesis of a Pauli channel is formulated via the following constrained minimization problem of a probability distribution $\{p_{j}\}$ given a target unitary channel $\hat{\mathcal{U}}$ and a set of synthesized unitaries $\{\mathcal{U}_{j}\}$ as follows:
@@
When the set $\{\mathcal{U}_{j}\}$ is not appropriately generated, there does not exist any feasible solution for Eq. (S25). While numerical solvers such as Gurobi offer an automated way to relax the equation constraint, via the variable presolve, we find that this precomputation is done too aggressively so that the accuracy of the minimization is quite poor. Instead, in our work, we have done a logarithmic search of $g_{1}$, starting from machine precision of $g_{1}=10^{-16}$. We have defined the error crafting to be successful if we find feasible solution such that (1) $\min\frac{1}{2}\|\hat{\mathcal{U}}-\sum_{j}p_{j}\mathcal{U}_{j}\|_{\diamond}\leq(c+1)^{2}\epsilon^{2},$ and (2) $g_{1}\leq 10^{-12}$ are both satisfied.
As we discussed in Sec. S2, the minimization problem boils down to a linear programming problem under the magic basis. The minimization problem equivalent to Eq. (S25) is
@@
where $r^{(j)}=(r_{1}^{(j)},r_{2}^{(j)},r_{3}^{(j)},r_{4}^{(j)})^{T}\in\mathbb{R}^{4}$ corresponds to the magic basis representation of $\mathcal{U}_{j}\circ\hat{\mathcal{U}}^{-1}.$ After relaxing the equality constraint, the minimization problem is given as
[…]
where $g^{\prime}_{1}$ is the tolerance factor in this representation. We find that accuracy is much enhanced when the minimization (and relaxation) is done under this representation. In practice, we solve Eq. (S28) using LP implemented in Gurobi or SciPy, and check if the conditions are satisfied.
S3.2 Crafting depolarizing channel
Here, we provide numerical details regarding the error crafting of depolarizing channels. In similar to Sec. S3.1, the minimization problem is defined as
@@
${\rm minimize}\ \frac{1}{2}\left\|\hat{\mathcal{U}}-\sum_{j}p_{j}\mathcal{U}_{j}\right\|_{\diamond}\ \ \ {\rm subject\ to}\ \ \sum_{a}\chi_{aa}=1,\max_{a^{\prime}\neq b^{\prime}}\{|\chi_{a^{\prime}a^{\prime}}-\chi_{b^{\prime}b^{\prime}}|\}\leq g_{2},\sum_{a\neq b}|\chi_{ab}|\leq g_{1},$ (S30)
@@
where $a\in\{{\rm I},{\rm X},{\rm Y},{\rm Z}\}$ and $a^{\prime},b^{\prime}\in\{{\rm X},{\rm Y},{\rm Z}\}.$ We regard the error crafting to be successful if (1) $\min\frac{1}{2}\|\hat{\mathcal{U}}-\sum_{j}p_{j}\mathcal{U}_{j}\|_{\diamond}\leq(c+1)^{2}\epsilon^{2}$, (2) $g_{1}\leq 10^{-12}$, and (3) $g_{2}\leq 0.01(\sum_{a^{\prime}\in\{X,Y,Z\}}\chi_{a^{\prime}a^{\prime}})\sim 0.01\epsilon^{2}$ are all satisfied. For completeness, we mention that the minimization problem can also be expressed in the magic basis representation as
[…]
where ${\mathbf{e}}^{(i)}$ is a vector whose element is 0 except the $i$th element. After relaxing the equality into inequality, we now have
minimize ${\rm minimize}\ 1-\sum_{j}p_{j}\left(r^{(j)}_{1}\right)^{2}$
${\rm subject\ to}\ \ \sum_{j}p_{j}=1,p_{j}\geq 0,\ \ \max_{k\neq k^{\prime}}\left\{\left|\sum_{j}p_{j}(r_{k}^{(j)}r_{k}^{(j)}-r_{k^{\prime}}^{(j)}r_{k^{\prime}}^{(j)})\right|\right\}\leq g^{\prime}_{2},\ \ \sum_{k\neq k^{\prime}}\left|\sum_{j}p_{j}r^{(j)}_{k}r^{(j)}_{k^{\prime}}\right|\leq g^{\prime}_{1},$ (S32)
where $g^{\prime}_{2}$ is the tolerance factor for the homogeneity of the Pauli error rates.
@@
The results of error-crafted synthesis are shown in Fig. S1. We observe the validity of the nine-shift unitary from the bounded behavior in Fig. S1(a). As in the case of the Pauli constraint, we also observe a boost in the accuracy by expanding the size of the basis set by a factor of $R$. From Figs. S1(b), (c), and (d), we further find that both increasing $c$ and $R$ contribute to the suppression of the failure probability of the error crafting.
@@
Figure S1: (a) Diamond distance $d_{\diamond}(\hat{\mathcal{U}},\mathcal{U}_{\rm synth})$ between the target single-qubit Haar random unitary $\hat{\mathcal{U}}$ and synthesized channel $\mathcal{U}_{\rm synth}\coloneqq\sum_{j}p_{j}\mathcal{U}_{j}$, with the remnant error constrained to the single-qubit depolarizing channel. The black real line indicates the results from unitary synthesis by the Ross-Selinger algorithm Ross and Selinger (2016), while the filled dots are results from error-crafted synthesis with shift factors of $c=2,3,5,7.$ When we expand the basis set size by a factor of $R=3$, shown by unfilled triangles, the effective synthesis error is suppressed below the bound in Eq. (S72) and approaches the value of $\epsilon^{2}$. The shown data are mean over 200 random instances of target unitaries. (b) Failure rate of error-crafted synthesis, averaged over 50 instances. (c)(d) Suppression of $p_{\rm fail}$ with shift factor $c$ and $R$. The data is averaged over 10000 and 2000 instances, respectively.
S3.3 Crafting XY channel
We finally describe the minimization problem to craft XY channels. Namely, we wish to solve the following:
${\rm minimize}\ \frac{1}{2}\left\|\hat{\mathcal{U}}-\sum_{j}p_{j}\mathcal{U}_{j}\right\|_{\diamond}\ \ \ {\rm subject\ to}\ \ \chi_{XX}+\chi_{YY}=1,\chi_{\rm ZZ}=0,\chi_{a\neq b}=0,$ (S33)
where the set of shift unitary $\{\mathcal{V}_{j}^{(c\epsilon)}\}$ is generated according to the procedure shown in Sec. S5. To enhance the numerical stability, we have relaxed the equality constraint for $\chi_{ZZ}$ into inequality as
${\rm minimize}\ \frac{1}{2}\left\|\hat{\mathcal{U}}-\sum_{j}p_{j}\mathcal{U}_{j}\right\|_{\diamond}\ \ \ {\rm subject\ to}\ \ \sum_{a}\chi_{a}=1,\max_{a^{\prime}\neq b^{\prime}}\{|\chi_{ZZ}|\}\leq g_{2},\sum_{a\neq b}|\chi_{a\neq b}|\leq g_{1},$ (S34)
where tolerance factors $g_{1}$ and $g_{2}$ are chosen in similar to those for depolarizing channels. Using the magic basis, the minimization problem is expressed as
minimize ${\rm minimize}\ 1-\sum_{j}p_{j}\left(r^{(j)}_{1}\right)^{2}$
${\rm subject\ to}\ \ \sum_{j}p_{j}=1,p_{j}\geq 0,\ \ \sum_{j}p_{j}r_{2}^{(j)}r_{2}^{(j)}=0,\ \ \sum_{j}p_{j}r^{(j)}r^{(j)T}{\rm\ is\ diagonal},$ (S35)
and after the relaxation of the equality constraint into inequality we have
minimize ${\rm minimize}\ 1-\sum_{j}p_{j}\left(r^{(j)}_{1}\right)^{2}$
${\rm subject\ to}\ \ \sum_{j}p_{j}=1,p_{j}\geq 0,\ \ \left|\sum_{j}p_{j}r_{2}^{(j)}r_{2}^{(j)}\right|\leq g^{\prime}_{2},\ \ \sum_{k\neq k^{\prime}}\left|\sum_{j}p_{j}r^{(j)}_{k}r^{(j)}_{k^{\prime}}\right|\leq g^{\prime}_{1},$ (S36)
where $g^{\prime}_{2}$ is the tolerance factor for suppressing the Z component of the Pauli channel and $g^{\prime}_{1}$ is for non-unital components.
@@
S4 Numerical details on error crafting using CPTP maps
As discussed in the main text, the error of Pauli rotation can be suppressed significantly by applying the error crafting for CPTP maps that represents the channel of synthesized gates accompanied by error correction circuit as shown in Fig. S2. In similar to the cases in Sec. S3, the minimization problem is defined as
$\frac{1}{2}\left\|\hat{\mathcal{U}}-\sum_{j}p_{j}\Lambda_{j}\right\|_{\diamond}\ \ \ {\rm subject\ to}\ \ \sum_{j}p_{j}\Lambda_{j}\circ\hat{\mathcal{U}}^{-1}(\cdot)=\sum_{a}\chi_{aa}P_{a}\cdot P_{a},\chi_{aa}\geq 0,$ (S37)
where $\{\Lambda_{j}\}$ is the set of implementable CPTP maps. Concretely for the case of Pauli rotation, with the target unitary and synthesized unitary channels expressed as $\hat{\mathcal{U}}$ and $\mathcal{U}_{j}$, respectively, we can express the effective CPTP map $\Lambda_{j}$ and remnant error $\mathcal{E}_{{\rm rem},j}$ obtained from logical error correction as
@@
Without loss of generality, we consider the single-qubit Pauli Z rotation in the following. First, note that, for the case of Pauli rotations, the goal of finding an appropriate set of CPTP maps and their weights $\{(p_{j},\Lambda_{j})\}_{j}$ is equivalent to finding an appropriate set of $\{p_{j},\mathcal{E}_{{\rm rem},j}\}_{j}$. It is straightforward to see that the Pauli transfer matrix of a remnant error channel $\mathcal{E}_{{\rm rem},j}$ for a Pauli Z rotation, denoted as $\Gamma_{\mathcal{E}_{{\rm rem},j}}$, can be expressed as
@@
where $0\leq\mu_{j}\leq 1$ and $\Gamma_{\mathcal{E}_{{\rm rem},j}}\geq 0.$ This motivates us to find a pair of remnant error channels $\mathcal{E}_{{\rm rem},+}$ and $\mathcal{E}_{{\rm rem},-}$ which satisfies (i) $\nu_{+}>0,\nu_{-}<0$, (ii) $\mu_{\pm}=O(\epsilon^{3})$, which follows from the geometric interpretation described in the main text. Once such a pair is found, we take the mixture of them as
@@
which is a purely Z error channel with an error rate of $\tilde{\mu}=\frac{\nu_{-}}{\nu_{+}+\nu_{-}}\mu_{+}+\frac{\nu_{+}}{\nu_{+}+\nu_{-}}\mu_{-}=O(\epsilon^{3}).$
We note that the current search algorithm is based on a brute-force search, and hence the computational complexity is expected to scale as $O(K/\epsilon)$ where $K$ is the number of shift unitaries. We leave an open question of whether there is any bound on complexity to search the pair as in the above discussion. In the current implementation, we generate around 3000 unitaries to reach a remnant error of $10^{-6}$ and around 20000 unitaries to reach $10^{-9},$ with each unitary synthesis done in milliseconds using a direct search algorithm proposed in Ref. Morisaki et al. (2024). Note that each unitary synthesis is independent, and hence is embarrassingly parallelizable.
S5 Generation of shift unitaries with fixed diamond distance
@@
S6 White noise approximation under coherent errors
@@
The white noise has been argued to be realized in experiments, in particular when the unitary gates are chosen at random Arute et al. (2019). In fact, one can prove that, even if the qubit connectivity is only linear and the circuit constitutes a brick-wall structure with random Haar unitary gates for two qubits, the effective noise converges to the white noise under unital noise Dalzell et al. (2024); Deshpande et al. (2022). However, these arguments mainly consider incoherent error, and how coherent error can affect the white-noise approximation is not fully examined. Given that the synthesis error is inevitable in the early FTQC regime, it is crucial to investigate the effect of coherent errors on the white-noise approximation.
@@
by choosing the rescaling factor as $R=(s/u)^{L}$ (Theorem S2 of Tsubouchi et al. (2024)). Here, $u$ and $s$ are noise-dependent constants called unitarity and average noise strength, respectively, which can be represented as
@@
We can observe a similar scaling numerically. Before discussing the effect of the coherent noise, let us briefly summarize the numerical argument for incoherent noise, especially the local depolarizing noise $\mathcal{E}=\bigcirc_{i=1}^{n}\mathcal{E}^{(i)}_{\mathrm{dep}}$, provided in Ref. Tsubouchi et al. (2023). We analyze the convergence to white noise by examining the Pauli transfer matrix (PTM) of the effective noise $\mathcal{E}^{\prime}$, denoted hereafter as $\Gamma$. We define $\Gamma$ by $\Gamma_{ij}={\rm Tr}[P_{i}\mathcal{E}^{\prime}(P_{j})]/2^{n}$ for $P_{i},P_{j}\in\{I,X,Y,Z\}^{\otimes n}$ for $n$-qubit system. Note that PTM is diagonal if $\mathcal{E}^{\prime}$ is a Pauli channel, and furthermore there exists $0\leq w\leq 1$ such that $\Gamma_{ii}=w~{}(i\neq 0)$ if $\mathcal{E}^{\prime}$ is the white noise. In this regard, it is useful to compute the singular values of the PTM in order to investigate the closeness to the white noise. We in particular compute $\Gamma=V\Lambda W^{\dagger}$ with $\Lambda={\rm diag}\{\lambda_{i}\}_{i=1}^{4^{n}}$, and define the damping factor $k_{i}$ as follows,
@@
It can be proven that, when each unitary layer consists of unitary 2-design, the damping factor converges to $k_{\rm mean}=\frac{3n}{4}\frac{4^{n}}{4^{n}-1}$, and numerical analysis supports that this also holds for weaker randomness, e.g., 2-local random circuits with linear connectivity, when the circuit depth is sufficiently deep. In other words, the mean deviation $\mathbb{E}_{i}[|k_{i}-k_{\rm mean}|]$, as well as the deviation of extremals $|k_{\rm max(min)}-k_{\rm mean}|$, is suppressed as $O(1/\sqrt{L})$ Tsubouchi et al. (2023), which is consistent with the suppression of total variation distance in measurement probability distribution in the computational basis Dalzell et al. (2024).
Now we ask how the above behavior is affected by the presence of coherent errors. Here, for the sake of simplicity of the argument, we sample $U_{i}$ uniformly from the entire unitary group, i.e., take an $n$-qubit Haar random unitary, and consider a local noise as $\mathcal{E}_{l}=\bigcirc_{i=1}^{n}\mathcal{E}_{\rm coh}^{(i)}\circ\mathcal{E}_{\rm dep}^{(i)},$ where $\mathcal{E}_{\rm coh}^{(i)}$ is a random 1-qubit unitary such that $\|\mathcal{E}_{\rm coh}^{(i)}-{\rm Id}\|_{\diamond}=\epsilon_{\rm coh}.$ In practice, this can be realized by constraining the rotation angles under the axial decomposition of the unitary.
Interestingly, we find that the effect of coherent error on the magnitude of the white noise is quadratically suppressed, while that on the convergence rate is linear. In Fig. S3(a)(b) we can see that, although the fluctuation in the damping factor is expanded at the shallow-depth regime, it readily converges at the rate of $O(1/\sqrt{L})$, which is in similar to the depolarizing-only case Tsubouchi et al. (2023). We observe that the deviation $\max_{k}\{|k-k_{\rm mean,dep}|\}$ is lifted approximately by a factor of $(\epsilon_{\rm coh}+\epsilon_{\rm dep})/\epsilon_{\rm dep}$ where $\epsilon_{\rm dep}:=\|\mathcal{E}_{\rm dep}^{(i)}-{\rm Id}\|_{\diamond}$, implying the slow down of the white noise convergence. Meanwhile, as shown in Fig. S3(c), we notice that the converged effective channel is less affected, in a sense that $|k_{\rm mean,tot}-k_{\rm mean,dep}|\propto\epsilon^{2}.$ Phenomenologically, we can understand that the effect of coherent noise is self-twirled by the circuit such that the diamond norm is suppressed quadratically as in the random compilation scheme Odake et al. (2024). These results indicate that, although the presence of coherent error does not severely increase the sampling overhead to perform error countermeasures such as the rescaling technique, the adequacy of the approximation is degraded and hence desirable to avoid them by error-crafted synthesis.
@@
S7 Theoretical guarantees for error crafted synthesis
S7.1 Crafting Pauli channel (Proof of Theorem 1)
@@
[…]
@@
where $\hat{A}_{6}$ is the matrix defined in the same way as Eq. (S53) for $\hat{\mathcal{U}}_{j}\circ\hat{\mathcal{U}}^{-1}$, and we use $\hat{A}_{6}^{-1}\hat{m}^{(7)}=(D\hat{A}_{6})^{-1}D\hat{m}^{(7)}$ in the last equality. Note that $\hat{A}_{6}$ also allows us to determine $\hat{m}^{(i)}~{}(i=1,...,6)$. If we can show that there exists an upper bound $u(c,\epsilon)$ of $\left\|Dm^{(j)}-D\hat{m}^{(j)}\right\|_{2}$ satisfying
@@
where we use the Cauchy-Schwartz inequality $\left(\sum_{i=1}^{3}\left(\hat{v}_{i}^{2}-\frac{\sin^{2}\theta}{\sin^{2}\hat{\theta}}v_{i}^{2}\right)^{2}\right)\left(\sum_{i=1}^{3}\left(\frac{1}{\sqrt{3}}\right)^{2}\right)\geq\left(\sum_{i=1}^{3}\frac{1}{\sqrt{3}}\left(\hat{v}_{i}^{2}-\frac{\sin^{2}\theta}{\sin^{2}\hat{\theta}}v_{i}^{2}\right)\right)^{2}$ in the last inequality. We proceed with the calculation as follows:
@@
S7.2 Crafting depolarizing channel
@@
[…]
@@
S8 Technical lemmas
S8.1 Uniform continuity of $g_{c}(\epsilon,t)$
@@
S8.2 Commutativity of $\lim$ and $\max$ for uniformly continuous functions